OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5

OpenStudy (anonymous):

(−10, −5) (−7, −2) (7, 2) (5, 10)

hartnn (hartnn):

isolate 'y' from the 2nd equation y-x=5 Add 5 on both sides.

hartnn (hartnn):

i mean add 'x' on both sides

OpenStudy (anonymous):

I add x to the 5? and x or y?

hartnn (hartnn):

like this : \(y-x = 5 \\ y-x+x = 5+x\) simplify that, y =... ?

OpenStudy (anonymous):

5?

hartnn (hartnn):

what about x ?

OpenStudy (anonymous):

10?

hartnn (hartnn):

no lol \(y-x+x = x+5\) -x + x will be 0 so you get \(y =x+5\) got this ?

OpenStudy (anonymous):

okay...

hartnn (hartnn):

now substitute this 'y' in the first equation, plug in y = x+5 in \(y^2+x^2 = 53\) what do u get ?

OpenStudy (anonymous):

@hartnn I honestly don't know...

hartnn (hartnn):